1) If the flat lamp on the outside of the IGRINS cal unit is staying on, take a flat image to reset the power state in the control software. Or press the power button inside the electronics cabinet.

2) If the auto-guide software is over-correcting and the source is oscillating around the slit, change the scaling factor in the offending direction.

3) If icsrun start-up routine fails to start and reports an error with the logging routine then reset permissions with:

igrins% sudo sh icsrun

4) When the Instrument Control Computer (ICC) has the screen freeze, look for an unresponsive script in the background:

command-option-escape

5) If you need to reinitialize the detector than follow the same steps as initialization in the Old Full Startup Procedures, except do not reload the JADE2 firmware.

6) Growling/Gurgling sounds from the coldhead may mean that the power cable has been compromised and needs to be replaced. More information available here.

7) If 'Go R' fails then first check that the TCS says 'On Target'. If the target is very bright then the centroid will fail and 'Go R' will fail. Moving the field manually, using the buttons above 'Go R' can sometimes fix this issue. Otherwise, please restart the control software.

8) When power was lost for 1 minute everything came back online except the IGRINS helium compressor. The breaker switch (B6) is located in room 107 on the first floor, next to the room where the glycol chiller is located.

9)Socket errors, ibunny.py not working, or SVC and DTP exposures not being taken?  First try restarting the ICC and following the Starting up IGRINS computers & data taking software.  If that doens't work try restarting the DSC and ICC and Starting up IGRINS computers & data taking software.  If that doesn't work try shutting down the DSC & ICC, wait 5 minutes, then start them up again.

If that still doesn't work you will need to fully reset IGRINS.  Shut down the detector control computers as discussed in Old Full Shutdown Procedures, then shut down DSC and ICC.  Unplug IGRINS power from the power socket in the dome floor.  Wait a minute.  Plug it back in.  Go up to the computer rack and turn off the power to the arc and flat lamps, and check the Lakeshores and turn the heaters on as described in Lakeshore Temperature and Heater Settings.  Push the power buttons on the detector control computers.  Close the computer rack up and head back to the control room. Turn DSC back on.  Turn ICC back on.  Follow the Old Full Startup Procedures for reactivating the detector computers and starting up the software on the ICC.  Hopefully this should work.

10) When trying to set up the guiding and clicking "Move A" or "Move B" does not work and you see  "AG Failure" error in the DTP log and the SVC does not seem to display new images:

Close the DTP, SVC, and Housekeeping packages.  Reload ibunny.py.  Rerun "bash icsrun".  If this doesn't solve the problem, close and reload the packages a second time.  That should fix it.

11) SVC package keeps saying "Finder is not found" even though your target is in the A or B box when guiding:

Resize the guider box slightly.  Try nudging the telescope west (or east) an arcsecond or two.

13) Clicking "Go R" in the SVC package only aligns the target parallel to the slit, and does not movie it perpendicular to the slit.

Try clicking your target and hitting "Go R" again, and it should work.  Cause of this glitch is unknown. 

  • No labels